One pipe can fill a tank in 20 minutes. A second pipe can fill the tank in 30 minutes. If the tank is empty, how long would be required for the two pipes operating together to fill it?

I need the rate of work (R), the time of work (T) , and the work done (W)
the equation is RT = W.
It would help if I am provided a detailed explanation

Answer by stanbon(75887)

One pipe can fill a tank in 20 minutes. A second pipe can fill the tank in 30 minutes. If the tank is empty, how long would be required for the two pipes operating together to fill it?
I need the rate of work (R), the time of work (T) , and the work done (W)
the equation is RT = W.
It would help if I am provided a detailed explanation
0 solutions
---------------------
1st Pipe DATA:
time = 20 min/job ; rate + 1/20 job/min
----------------------
2nd Pipe DATA;
time = 30 min/job ; rate = 1/30 job/min
----------------------
Together DATA:
time = x min/job ; rate = 1/x job/min
---------------------------------------
Equation:
rate + rate = together rate
1/20 + 1/30 = 1/x
---
Multiply thru by x to get:
work + work = job done
x/20 + x/30 = 1
Multiply thru by 60:
3x + 2x = 60
5x = 60
x = 12 min (time needed to complete the job when both pipes are operating)
=======================
Cheers,
Stan H.